# MEG: Medical Knowledge-Augmented Large Language Models for Question Answering

### v2

@@ Abstract
Question answering is a natural language understanding task that involves reasoning over both explicit context and unstated, relevant domain knowledge. Large language models (LLMs), which underpin most contemporary question answering systems, struggle to induce how concepts relate in specialized domains such as medicine. Existing medical LLMs are also costly to train. In this work, we present MEG, a parameter-efficient approach for medical knowledge-augmented LLMs. MEG uses a lightweight mapping network to integrate graph embeddings into the LLM, enabling it to leverage external knowledge in a cost-effective way. We evaluate our method on four popular medical multiple-choice datasets and show that LLMs greatly benefit from the factual grounding provided by knowledge graph embeddings. MEG attains an average of +10.2% accuracy over the Mistral-Instruct baseline, and +6.7% over specialized models like BioMistral. We also show results based on Llama-3. Finally, we show that MEG’s performance remains robust to the choice of graph encoder.
@@ 1 Introduction
To improve reliability and reduce computational costs, researchers have experimented with training from mixtures of corpora and knowledge bases (Pan et al., 2023, 2024). Knowledge Graphs (KGs), such as the Unified Medical Language System (UMLS) Bodenreider (2004), are structured knowledge bases that explicitly store rich factual knowledge. KGs are good at capturing the nuances of complex data and can provide complementary information to LLMs, especially useful for tasks requiring structured understanding. The potential of knowledge-augmented LLMs [^1] In this work, we define a knowledge-augmented LLM as an LLM enhanced with KG embeddings (KGEs). KGEs are dense vector representations of graph entities Ju et al. (2024). Therefore, we also refer to knowledge-augmented LLMs as KGE-augmented LLMs throughout the paper. outlines an interesting research paradigm that can alleviate current challenges of LLMs, and reduce the need of training ever-larger models Hooker (2024). However, how to effectively model interactions between LLMs and KGs remains an open question.
@@
In this work, we introduce MEG, a parameter-efficient approach to MEdical knowledGe-augmented LLMs for question answering (QA). We design a lightweight mapping network to unidirectionally translate KG embeddings into the LLM’s vector space. This enables the LLM to interpret the new input embeddings, which, in turn, further conditions its response generation. We use Mistral-Instruct (7B) Jiang et al. (2023) as our base LLM and report results with our best setup: a KG encoder based on GraphSAGE Hamilton et al. (2017) combined with a simple MLP as mapping network. We also provide results with the recently released Llama-3-Instruct (8B) Dubey et al. (2024) as base LLM.
In sum, our contributions are as follows: i) We introduce MEG, a novel approach to knowledge-augmented LLMs based on KGEs. ii) We conduct extensive evaluation on the four popular multiple-choice QA datasets from the MultiMedQA Singhal et al. (2023a) clinical benchmark, and demonstrate the effectiveness of integrating pretrained KGEs into LLMs for medical question answering. Specifically, MEG surpasses strong LLM baselines like BioMistral-7B Labrak et al. (2024) or MediTron-7B Chen et al. (2023), which have followed a costly continued pretraining of the base LLMs on curated biomedical data. iii) We provide insights into the inner workings of MEG, examining the contributions of each module and comparing embedding spaces. We intuitively explain the shifts in the LLM’s representations that drive MEG’s stronger performance. iv) We publicly release the code, trained KGEs and model checkpoints at github.com/lautel/MEG.
@@ 2 Related Work
Current state-of-the-art (SOTA) in medical QA benchmarks like MedQA Jin et al. (2020), PubMedQA Jin et al. (2019) or MedMCQA Pal et al. (2022) belongs to close-sourced models of unknown size like Med-Gemini Saab et al. (2024), Med-PaLM2 Singhal et al. (2023b) or GPT-4 Nori et al. (2023). Popular open-source LLMs in biomedicine include MedAlpaca Han et al. (2023) and PMC-LLaMA Wu et al. (2023) based on Llama Touvron et al. (2023a), MediTron Chen et al. (2023) based on Llama-2 Touvron et al. (2023b), or BioMistral Labrak et al. (2024) based on Mistral-Instruct Jiang et al. (2023). These models continue pretraining the base general-purpose models on curated medical corpora. More recently, Kim et al. (2024) present the Meerkat models trained with chain-of-thought Wei et al. (2024) synthetic data. Meerkat-7B outperforms the previous best 7B models across several medical benchmarks. However, it takes eight 80G A100 GPUs and 1.5 days to complete training. In contrast, our approach is the first to leverage pretrained medical KGEs and can be trained on four A10G GPUs within a few hours (see § 5 for details).
@@
Bringing together LLMs and KGs is an active line of research that has gained increasing attention from both academia and industry (Pan et al., 2023, 2024). Among numerous efforts in this area, Zhang et al. (2019); Yasunaga et al. (2022); Tang et al. (2024); Zhu et al. (2023), to name a few, propose different methods for combining text and graphs during pretraining. Parallel to these lines of work, Sarmah et al. (2024); Edge et al. (2024); Hu et al. (2024); Mavromatis and Karypis (2024) approach the integration of LLMs and KGs through retrieval-augmented generation (RAG) Lewis et al. (2020). However, the deployment of such knowledge-augmented LLMs for medical QA remains understudied. Our work fills this gap and presents a novel approach to medical knowledge-augmented LLMs based on KGEs. We note that MEG may resemble a sort of RAG system, where an LLM leverages knowledge from an external database of KGEs. In this case, the grounding module would act as retrieval module, fetching appropriate KGEs that ground text information in KG entities as part of a prompt.
@@ 3 Problem Formulation
We augment an LLM with KG embeddings to answer medical questions drawing on factual knowledge from the KG. We rely on a large KG in the target domain, namely UMLS Bodenreider (2004). Our proposed approach, MEG, consists of four key components:
@@
Figure 1 depicts the full pipeline of MEG. We carefully investigate the design of these components and how they interact with each other (§ 4.1). To attain the best accuracy on downstream tasks, we conduct a two-phase training (§ 4.2).
@@
Figure 1: MEG leverages a pretrained KG encoder and an LLM. During an initial phase of training, MEG learns a mapping network to convert relevant graph features (KGEs) retrieved by the grounding module into token embeddings. During downstream fine-tuning, only the LLM’s weights are updated, keeping the LLM’s embedding layer and mapping network frozen. At inference, the LLM takes the text and the mapped KGEs as input and generates a response.
@@ 4.1 MEG
MEG combines a pretrained KG encoder and a pretrained LLM by means of an intermediate mapping network (see Figure 1). The KG encoder, which is trained separately on a large medical KG [^3] Specifically, we use UMLS Bodenreider (2004), a widely-used KG in biomedicine with $\sim$300K nodes (entities) and one million edges in total., provides graph embeddings that are directly fed to the mapping network. Then, the LLM uses the text content and mapped KG embeddings as input to generate an answer.
@@
The goal is to learn the mapping $f_{\text{k}}$ that transforms $X$ to the LLM’s vector space, while preserving its semantic meaning and structural information. Rather than minimizing the sum of squared differences between $f_{\text{k}}(X)$ and $Y$, we aim at positioning each $x_{i}$ in the neighborhood of its counterpart in $Y$. Pursuing an exact matching of space distributions, such as through a Procrustes transformation Schönemann (1966); Gower (1975), would disregard the structural knowledge encoded in $X$.
To achieve this, we design an architecture similar to Xu et al. (2018) with two mappings $f_{\text{k}}:X\rightarrow Y$ and $g_{\text{k}}:Y\rightarrow X$, as illustrated in Figure 2. We construct an instruction dataset with labels from UMLS’s entities to teach the LLM to interpret the transformed graph embedding $f_{\text{k}}(x_{i})$. Figure 2 shows an example of an instruction, where the placeholder {kg_embedding} is replaced by $f_{\text{k}}(x_{i})$. We train the full network jointly with the LLM [^4] We conducted experiments by training the mapping network and LLM separately on UMLS. However, this approach resulted in worse performance on the downstream tasks tested.. Our loss function consists of three parts: a standard next-token prediction objective (cross-entropy loss $\mathcal{L_{\text{ce}}}$), and a sum of a contrastive loss and a back-translation loss to optimize the mapping network. Specifically,
@@
The grounding module takes textual data $W$ as input and links entity mentions in $W$ to their corresponding nodes in the knowledge graph $G$, generating a set of KG embeddings for the LLM to use. A critical step in this process is medical entity disambiguation Vretinaris et al. (2021); Lu et al. (2024), which involves detecting named entities in $W$ and linking them to their unique counterparts in $G$. Since an entity can be referred to in multiple ways, for instance ‘heart attack’ and ‘myocardial infarction’, this step standardizes variations by linking them to a unique identifier in $G$ (following with the previous example, both mentions correspond to Concept Unique Identifier (CUI) ‘C0155626’ in UMLS). This grounding ensures retrieval of relevant information for each example. We use the entity linker presented in Neumann et al. (2019) [^6] We use the last version of scispaCy (v2.5.0), which supports linking to UMLS and has near 3M unique concepts., which covers 99% of the concepts mentioned in the MedMentions dataset Mohan and Li (2019) and 86% of the concepts mentioned in the MedQA dataset Jin et al. (2020). These two datasets provide ground truth UMLS annotations.
@@ 4.2 Training
We aim to train MEG to achieve competent results on medical question answering benchmarks while minimizing computational cost. To do this, we conceive a two-phase training strategy with a minimal part of the model’s parameters updated.
@@
We first learn the optimal transformation $f_{\text{k}}:X\rightarrow Y$ so that the mapped KG embeddings retain relevant information from the KG and can be effectively used by the LLM. As explained in § 4.1, we create an instruction dataset to guide the LLM in learning the relationship between its original representation of medical entities and their mapped graph embeddings. The train set contains 297,927 examples, following the same template shown in Figure 2 for every entity label in UMLS. [^7] We investigate whether data augmentation at this stage could lead to more accurate results in downstream tasks. We augment the initial $\sim$300K examples by creating new instructions with multiple entities, e.g.‘Explain to me these medical concepts: […]’, to better match the setting from downstream tasks, which often include several entities per sample. This process doubles the dataset size, with an augmentation that normally distributes the number of entities per instruction between 2 and 10. Results are within ±$\pm 0.2$ accuracy in MedQA compared to training without the augmented data. Due to the extra computational costs and minor (if any) gains, we did not explore this option further. We train for one epoch jointly the mapping network and the LLM to minimize the objective from Eq. 3.
Phase II: Downstream task
Given a medical multiple choice QA dataset, we fine-tune MEG to answer the input question based on the textual content and information leveraged from the mapped KG embeddings. We format the input prompts $W$ as follows. For each example in the dataset, we concatenate the context (if any), question and candidate answer set following the pseudo-code shown in Figure 3. The placeholder {kg_embedding} is replaced with $N$ KGEs transformed by the mapping network888 In both training phases, we investigate the effect of injecting the mapped KGEs at the last layer of the LLM instead of after the embedding layer. These early experiments revealed little to no impact on zero-shot downstream accuracy, but slightly worsen the fine-tuned accuracy as measured on the validation set of MedQA with three random initialization seeds. This finding suggests that the LLM benefits from attending the external KGEs during fine-tuning, enabling more contextualized representations of these embeddings.. We assume $f_{\text{k}}$ is learned in phase I, so we keep the mapping network weights frozen and disconnect the backward network $g_{\text{k}}$, which merely served to regulate the learning of $f_{\text{k}}$ and prevent degenerated transformation. Similarly, the LLM’s embedding layer is also frozen.
@@ 5 Experimental Details
Following previous research on medical LLMs, we evaluate MEG on four well-known medical benchmarks that require extensive background knowledge. The first one is MedQA-USMLE (MedQA) Jin et al. (2020), which consists of 10,178 train questions and 1,273 test questions, formatted with four choices each. The content was originally curated by experts from the US Medical License Exam. The second benchmark, PubMedQA Jin et al. (2019), was collected from PubMed abstracts and includes 1,000 expert labeled question-answer pairs. The task is to produce a yes/no/maybe answer based on the question and an abstract as context. As previously done by others Singhal et al. (2023a); Chen et al. (2023); Labrak et al. (2024), we use 500 random [^9] We split the data following a similar distribution of answers between train and test splits. samples for evaluation. The remaining 500 samples, though limited in size, serve as our only source of training data. We exclude the 211k artificially labeled yes/no samples provided by Jin et al. (2019) to avoid bias towards these two options. The third benchmark, MedMCQA Pal et al. (2022), contains 179,722101010We detect 3,100 duplicate questions in the train split, which we remove. train questions from Indian medical entrance exams. Due to the unavailability of answer keys for the test set, we follow others Wu et al. (2023); Tu et al. (2023); Labrak et al. (2024) and report results on the validation set (4,183 questions). Lastly, MMLU-Medical Singhal et al. (2023a) includes 1,089 questions, each with four options, across six medical and biology-related categories drawn from Hendrycks et al. (2021). Since this dataset only provides test data, we evaluate the generalization performance of MEG fine-tuned on MedMCQA as in Chen et al. (2023). Thus, results on MMLU-Medical report out-of-distribution inference.
@@
During phase I of training, described in § 4.2, we randomly initialize the mapping network and load the pretrained weights of the LLM. We fully train the mapping network and perform low-rank adaptation (LoRA, Hu et al. (2022)) fine-tuning on every linear layer of the LLM, while the remaining parameters are frozen. This parameter-efficient tuning approach allows to learn the equivalent of 2% of the model’s parameters. Our full architecture results on 216M trainable parameters for MEG-Mistral. After fine-tuning, we merge the LLM’s updated parameters with the base model. Since the input prompt has a fixed size (see Figure 2), we use a reduced sequence length (124) to optimize computational efficiency. We train for one epoch with gradient accumulation over 8 steps to achieve an effective batch size of 128. We employ a cosine learning rate scheduler with learning rate of $51e-$ 5, warmup ratio of 3% and no weight decay. We use mixed precision (bfloat16) and FlashAttention2 Dao (2023) to optimize memory usage and speed up computations on the LLM. Training takes 4h on 4 NVIDIA A10G GPUS using DeepSpeed [^11] https://www.deepspeed.ai/ for distributed training.
After phase I, the models are evaluated in a zero-shot setup on downstream tasks, already able to use information from graph embeddings. To fine-tune on a specific task, we perform phase II of training. To allow batching, the number of KGEs injected to the LLM is fixed across samples121212 The average number of ground entities per instance varies across datasets according to the median number of ground entities. We set $N=20$ in MedQA, PubMedQA and ‘professional medicine’ in MMLU-Medical; $N=3$ in MedMCQA and $N=2$ in the remaining categories from MMLU-Medical.. If the grounding module retrieves more KGEs, we randomly select $N$. Otherwise, we add zero-padding. We assume the mapping network is learned in phase I, so we freeze its weights and disconnect the backward network $g_{\text{k}}$. We also freeze the LLM’s embedding layer. This approach reduces the computational complexity and speeds up fine-tuning with a total of 83M trainable parameters in both MEG-Mistral and MEG-Llama. Fine-tuning is done for 3 epochs, with a sequence length of 400 (500 for PubMedQA), learning rate of $41e-$ 4 and effective batch size of 32. The remaining hyperparameters have the same value as in phase I. We did not optimize hyperparameters for Llama-3-Instruct.
6 Results
We evaluate accuracy on four medical multiple-choice question datasets in three variants of MEG: MEG-Mistral1 and MEG-Mistral3, based on Mistral-7B-Instruct-v0.1 and -v0.3, respectively; and MEG-Llama, based on Llama-3-Instruct. We report average accuracy and standard deviation across three random seeds. Our results in Tables 1 and 2 reveal consistent average improvement across datasets compared to baselines.
In-prompt graph triples provide useful information

### v3

@@ Abstract
Question answering is a natural language understanding task that involves reasoning over both explicit context, and unstated relevant domain knowledge. Despite the high cost of training, large language models (LLMs)—the backbone of most modern question-answering systems—still struggle to reliably capture the nuanced relationships between concepts that are crucial for reasoning in specialized fields like medicine. In this work, we present MEG, a parameter-efficient approach for medical knowledge-augmented LLMs. MEG uses a lightweight mapping network to incorporate knowledge graph embeddings into the LLM, enabling it to leverage external knowledge in a cost-effective way. We evaluate our method on four popular medical multiple-choice datasets and show that LLMs i) can effectively interpret knowledge graph embeddings and ii) gain significant advantages from the factual grounding these embeddings provide. MEG attains an average of +6.7% and +9.9% accuracy over specialized models like BioMistral-7B and MediTron-7B, respectively. Finally, we show that MEG’s performance remains robust to the choice of graph encoder.
@@ 1 Introduction
To improve reliability and reduce computational costs, researchers have experimented with training from mixtures of corpora and knowledge bases (Pan et al., 2023, 2024). Knowledge Graphs (KGs), such as the Unified Medical Language System (UMLS) Bodenreider (2004), are structured knowledge bases that explicitly store rich factual knowledge. KGs are good at capturing the nuances of complex data and can provide complementary information to LLMs, especially useful for tasks requiring structured understanding. However, this complementary information is introduced as a new modality: graphs. Therefore, it extends beyond the plain text that LLMs can naturally interpret, requiring additional mechanisms to process and integrate it effectively. The potential of knowledge-augmented LLMs [^1] In this work, we define a knowledge-augmented LLM as an LLM enhanced with KG embeddings (KGEs). KGEs are dense vector representations of graph entities Ju et al. (2024). Therefore, we also refer to knowledge-augmented LLMs as KGE-augmented LLMs throughout the paper. outlines an interesting research paradigm that can alleviate current challenges of LLMs, and reduce the need of training ever-larger models Hooker (2024). However, how to effectively model interactions between LLMs and KGs remains an open question.
@@
In this work, we introduce MEG, a parameter-efficient approach to MEdical knowledGe-augmented LLMs for question answering (QA). MEG eliminates the need to train base LLMs for specialized domains by incorporating relevant KGEs at inference time. To achieve this, we design a lightweight mapping network that unidirectionally translates KGEs into the LLM’s vector space, allowing the model to interpret these embeddings, which, in turn, further conditions its response generation. Notably, KGEs are injected directly at token level, after the model’s embedding layer. We experiment with Mistral-Instruct (7B) Jiang et al. (2023) and Llama-3-Instruct (8B) Dubey et al. (2024) as our base models, and report results with our best setup for KGEs: a KG encoder based on GraphSAGE Hamilton et al. (2017) combined with a simple Multilayer Perceptron (MLP) as mapping network.
In sum, our contributions are as follows: [^2] Code and KGEs will be published upon acceptance. i) We introduce MEG, a novel approach to knowledge-augmented LLMs based on KGEs [^3] While we frame our approach in the medical domain, we would like to remark that it is domain independent given that it is built on any large KG. However, as we focused in this domain given the abundance of previous results and datasets, we refrain from making further generalized claims (in spite of there being strong evidences supporting them).. MEG is an efficient approach to domain specialization of models, as it updates only a small fraction of the LLM’s parameters while providing access to a large, reliable knowledge source (a KG). ii) We conduct extensive evaluation on the four popular multiple-choice QA datasets from the MultiMedQA Singhal et al. (2023a) clinical benchmark, and demonstrate the effectiveness of integrating pretrained KGEs into LLMs for medical question answering. Specifically, MEG surpasses well-establish biomedical LLM baselines like BioMistral-7B Labrak et al. (2024) or MediTron-7B Chen et al. (2023), which have followed a costly continued pretraining of the base LLMs on curated biomedical data. iii) We provide insights into MEG’s dual modality by examining the contributions of each module and comparing embedding spaces of KGEs and LLMs. We intuitively explain the shifts in the LLM’s representations that drive MEG’s stronger performance.
@@ 2 Related Work
Current state-of-the-art (SOTA) in medical QA benchmarks like MedQA Jin et al. (2020), PubMedQA Jin et al. (2019) or MedMCQA Pal et al. (2022) belongs to close-sourced models of unknown size like Med-Gemini Saab et al. (2024), Med-PaLM2 Singhal et al. (2023b) or GPT-4 Nori et al. (2023). Popular open-source LLMs in biomedicine include MedAlpaca Han et al. (2023) and PMC-LLaMA Wu et al. (2023) based on Llama Touvron et al. (2023a), MediTron Chen et al. (2023) based on Llama-2 Touvron et al. (2023b), or BioMistral Labrak et al. (2024) based on Mistral-Instruct Jiang et al. (2023). These models continue pretraining the base general-purpose models on curated medical corpora. More recently, Kim et al. (2024) present the Meerkat models trained with chain-of-thought Wei et al. (2024) synthetic data. Meerkat-7B outperforms the previous best 7B models across several medical benchmarks. However, it takes eight 80G A100 GPUs and 1.5 days to complete training. In contrast, our approach is among the first to leverage pretrained medical KGEs to condition the LLM’s generation and can be trained on four A10G GPUs within a few hours (see § 5 for details).
@@
Bringing together LLMs and KGs is an active line of research that has gained increasing attention from both academia and industry (Pan et al., 2023, 2024). Among numerous efforts in this area, Zhang et al. (2019); Yasunaga et al. (2022); Tang et al. (2024); Zhu et al. (2023), to name a few, propose different methods for combining text and graphs during pretraining. Parallel to these lines of work, Sarmah et al. (2024); Edge et al. (2024); Hu et al. (2024); Mavromatis and Karypis (2024) approach the integration of LLMs and KGs through retrieval-augmented generation (RAG) Lewis et al. (2020). However, the deployment of such knowledge-augmented LLMs for medical QA remains understudied. Our work fills this gap and presents a novel approach to medical knowledge-augmented LLMs based on KGEs. We note that MEG may resemble a sort of RAG system, where an LLM leverages knowledge from an external database of KGEs. However, our approach does not retrieve document chunks to include them in the textual prompt. Instead, MEG fetches KGEs of named entities from the user’s prompt and injects these embeddings into the LLM, enriching the query’s context and guiding answer generation. For a comprehensive benchmark of RAG in medical QA, we refer the reader to Xiong et al. (2024).
@@ 3 Problem Formulation
Our objective is to augment a general purpose LLM with specialized knowledge, so it can generate accurate answers without the need of fine-tuning to the new domain. We propose to use a large KG as the source of specialized knowledge. The reason for this is two-folded: a KG is a reliable source of factual knowledge and it captures structured relationship between concepts, that can potentially augment the LLM’s representation of such concepts. This design choice marks the multimodality of our solution: text generation will be conditioned on embedded representations of graph concepts.
Thus, our main task is to augment an LLM with KG embeddings to answer questions in a relevant domain. Our proposed approach, MEG, consists of four key components:
@@
Figure 1 depicts our full pipeline. We carefully investigate the design of these components and how they interact with each other (§ 4.1). The mapping network is jointly trained with the LLM to learn an appropriate transformation of KGE, and it is frozen in downstream applications (§ 4.2).
@@
Figure 1: MEG leverages a pretrained KG encoder and an LLM. During an initial phase of training, MEG learns a mapping network to convert relevant graph features (KGEs) retrieved by the grounding module into token embeddings. During downstream fine-tuning, only a subset of the LLM’s weights is updated, while the embedding layer and mapping network remain frozen. At inference, the LLM takes the text and the mapped KGEs as input and generates a response.
@@ 4.1 MEG
MEG combines a pretrained KG encoder and a pretrained LLM by means of an intermediate mapping network (see Figure 1). The KG encoder, which is trained separately on a large medical KG [^5] Specifically, we use UMLS Bodenreider (2004), a widely-used KG in biomedicine with $\sim$300K nodes (entities) and one million edges in total., provides graph embeddings that are directly fed to the mapping network. The mapping network outputs embeddings that are injected into the LLM after its embedding layer, influencing the LLM’s answer generation.
@@
The goal is to learn the mapping $f_{\text{k}}$ that transforms $X$ to the LLM’s vector space, while preserving its semantic meaning and structural information. Rather than minimizing the sum of squared differences between $f_{\text{k}}(X)$ and $Y$, we aim at positioning each $x_{i}$ in the neighborhood of its counterpart in $Y$. Pursuing an exact match of space distributions, such as through a Procrustes transformation Schönemann (1966); Gower (1975), would disregard the structural knowledge encoded in $X$.
To achieve this, we design an architecture similar to Xu et al. (2018) with two mappings $f_{\text{k}}:X\rightarrow Y$ and $g_{\text{k}}:Y\rightarrow X$, as illustrated in Figure 2. We construct an instruction dataset with labels from UMLS’s entities to teach the LLM to interpret the transformed graph embedding $f_{\text{k}}(x_{i})$. Figure 2 (top) shows an example of an instruction, where the placeholder {kg_embedding} is replaced by $f_{\text{k}}(x_{i})$ at token level. We train the full network jointly with the LLM [^6] We conducted experiments by training the mapping network and LLM separately on UMLS. However, this approach resulted in worse performance on the downstream tasks tested.. Our loss function consists of three parts: a standard next-token prediction objective (cross-entropy loss $\mathcal{L_{\text{ce}}}$), and a sum of a contrastive loss and a back-translation loss to optimize the mapping network. Specifically,
@@
The grounding module takes textual data $W$ as input and links mentioned entities in $W$ to their corresponding nodes in the knowledge graph $G$, generating the corresponding entity mentions’ KGEs for the LLM to use. These KG embeddings will enrich the question’s context, helping to guide the answer. A critical step in this process is medical entity disambiguation Vretinaris et al. (2021); Lu et al. (2024), which involves detecting named entities in $W$ and linking them to their unique counterparts in $G$. Since an entity can be referred to in multiple ways, for instance ‘heart attack’ and ‘myocardial infarction’, this step standardizes variations by linking them to a unique identifier in $G$ (following with the previous example, both mentions correspond to Concept Unique Identifier (CUI) ‘C0155626’ in UMLS). This grounding ensures retrieval of relevant information for each example. We use the entity linker presented in Neumann et al. (2019) [^8] We use the last version of scispaCy (v2.5.0), which supports linking to UMLS and has near 3M unique concepts., which covers 99% of the concepts mentioned in the MedMentions dataset Mohan and Li (2019) and 86% of the concepts mentioned in the MedQA dataset Jin et al. (2020). These two datasets provide ground truth UMLS annotations.
@@ 4.2 Training
We aim to train MEG to achieve competent results on medical question answering benchmarks while minimizing computational cost. To do this, We perform an initial training phase for embedding transfer learning.
@@
We first learn the optimal transformation $f_{\text{k}}:X\rightarrow Y$ so that the mapped KG embeddings retain relevant information from the KG and can be effectively used by the LLM. As explained in § 4.1, we create an instruction dataset from UMLS labels to guide the LLM in learning the relationship between its original representation of medical entities and their mapped graph embeddings. The train set contains 297,927 examples, following the same template shown in Figure 2 for every entity label in UMLS. [^9] We investigate whether data augmentation at this stage could lead to more accurate results in downstream tasks. We augment the initial $\sim$300K examples by creating new instructions with multiple entities, e.g.‘Explain to me these medical concepts: […]’, to better match the setting from downstream tasks, which often include several entities per sample. This process doubles the dataset size, with an augmentation that normally distributes the number of entities per instruction between 2 and 10. Results are within ±$\pm 0.2$ accuracy in MedQA compared to training without the augmented data. Due to the extra computational costs and minor (if any) gains, we did not explore this option further. We train for one epoch jointly the mapping network and the LLM to minimize the objective from Eq. 3. At this stage, MEG can leverage concepts from the KG to augment its domain knowledge and generate better, informed answers.
Fine-tuning on Downstream Tasks
Given a medical multiple choice QA dataset, we fine-tune MEG to answer the input question based on the textual content and information leveraged from the mapped KG embeddings. We format the input prompts $W$ as follows. For each example in the dataset, we concatenate the context (if any), question and candidate answer set following the pseudo-code shown in Figure 3.
At the token level, the placeholder embedding {kg_embedding} is replaced with $N$ transformed KGEs produced by the mapping network101010 In both training phases, we investigate the effect of injecting the mapped KGEs at the last layer of the LLM instead of after the embedding layer. These early experiments revealed little to no impact on zero-shot downstream accuracy, but slightly worsen the fine-tuned accuracy as measured on the validation set of MedQA with three random initialization seeds. This finding suggests that the LLM benefits from attending the external KGEs during fine-tuning, enabling more contextualized representations of these embeddings.. The mapping network weights remain frozen, and the backward network $g_{\text{k}}$ is effectively disabled, which previously served to regulate the learning of $f_{\text{k}}$ and prevent degenerated transformations. Similarly, the LLM’s embedding layer is also frozen. This approach reduces the computational complexity and speeds up fine-tuning.
@@ 5 Experimental Details
Following previous research on medical LLMs, we evaluate MEG on four well-known medical benchmarks that require extensive background knowledge. The first one is MedQA-USMLE (MedQA) Jin et al. (2020), which consists of 10,178 train questions and 1,273 test questions, formatted with four choices each. The content was originally curated by experts from the US Medical License Exam. The second benchmark, PubMedQA Jin et al. (2019), was collected from PubMed abstracts and includes 1,000 expert labeled question-answer pairs. The task is to produce a yes/no/maybe answer based on the question and an abstract as context. As previously done by others Singhal et al. (2023a); Chen et al. (2023); Labrak et al. (2024), we use 500 random [^11] We split the data following a similar distribution of answers between train and test splits. samples for evaluation. The remaining 500 samples, though limited in size, serve as our only source of training data. We exclude the 211k artificially labeled yes/no samples provided by Jin et al. (2019) to avoid bias towards these two options. The third benchmark, MedMCQA Pal et al. (2022), contains 179,722121212We detect 3,100 duplicate questions in the train split, which we remove. train questions from Indian medical entrance exams. Due to the unavailability of answer keys for the test set, we follow others Wu et al. (2023); Tu et al. (2023); Labrak et al. (2024) and report results on the validation set (4,183 questions). Lastly, MMLU-Medical Singhal et al. (2023a) includes 1,089 questions, each with four options, across six medical and biology-related categories drawn from Hendrycks et al. (2021). Since this dataset only provides test data, we evaluate the generalization performance of MEG fine-tuned on MedMCQA as in Chen et al. (2023). Thus, results on MMLU-Medical report out-of-distribution inference.
@@
During phase I of training, described in § 4.2, we randomly initialize the mapping network and load the pretrained weights of the LLM. We fully train the mapping network and perform low-rank adaptation (LoRA, Hu et al. (2022)) fine-tuning on every linear layer of the LLM (except for the embedding layer), while the remaining parameters are frozen. This parameter-efficient tuning approach allows to learn the equivalent of 2% of the model’s parameters. Our full architecture results on 216M trainable parameters.After training, we merge the LLM’s updated parameters with the base model. Training takes 4h on 4 NVIDIA A10G GPUS using DeepSpeed [^13] https://www.deepspeed.ai/ for distributed training. See Appendix A for further details.
After phase I, the models are evaluated in a zero-shot setup on downstream tasks, already able to use information from graph embeddings. To achieve optimal performance in downstream tasks, we fine-tune the models as in § 4.2. To allow batching, the number of KGEs injected to the LLM is fixed across samples141414 The average number of ground entities per instance varies across datasets according to the median number of ground entities. We set $N=20$ in MedQA, PubMedQA and ‘professional medicine’ in MMLU-Medical; $N=3$ in MedMCQA and $N=2$ in the remaining categories from MMLU-Medical.. If the grounding module retrieves more KGEs, we randomly select $N$. Otherwise, we add zero-padding. See Appendix A for training details.